A solution is made by dissolving 0.5g Na2CO3 in 10 mL of water, and that solution is then titrated with 0.5MHCl. Calculate the pH of the solution after adding 15 mL of the acid titrant. For H2CO3, pKa,1 = 6.35, pKa,2 = 10.33.

a. 12.55

b. 8.23

c. 6.18

d. 3.45
Answer #1

moles of Na2CO3 = 0.5 / 106 = 4.72 x 10^-3

moles of H+ = 15 x 0.5 / 1000 = 7.5 x 10^-3

CO32-   +   H+    -------------------> HCO3-

0.00472      0.0075                          0

0                 0.00278                     0.00472

HCO3-   + H+ -----------------> H2CO3

0.00472     0.00278                     0

0.00194                                     0.00278

pH = pKa1 + log [HCO3- / H2CO3]

pH = 6.35 + log (0.00194 / 0.00278)

pH = 6.19

answer : C) 6.18